OpenStudy (emilyjones284):

Write a system of two equations in two variables to solve the problem. A boat traveled 63 miles in 3 hours. Return trip took 7 hours going upstream. Find the speed of the boat in still water and the speed of the current.

OpenStudy (mhchen):

Model downstream rate = b+c and upstream rate = b-c 63 = (b + c) * 3 63 = (b - c) * 7 Reference: https://www.algebra.com/algebra/homework/equations/Equations.faq.question.206439.html

OpenStudy (emilyjones284):

Thank you! Could I also write it as x+3y=63 and x+7y=63?

OpenStudy (emilyjones284):

I just don't understand how I would solve a system of questions from the ones you gave me

OpenStudy (francesca_marie):

this might help http://www.themathpage.com/alg/word-problems2-2.htm

OpenStudy (emilyjones284):

Or actually 3x+3y=63 and 7x-7y=63?

OpenStudy (mhchen):

b is boat speed, c is water speed

OpenStudy (emilyjones284):

Would the equations that I said work?

OpenStudy (emilyjones284):

I got 10.5 for x and y

OpenStudy (mhchen):

Yes, your equations are the same :D